A program takes 20 seconds to run on 1 processor and 5 seconds to run on 4 processors. Estimate Tseq. What is the maximum speedup?
Answer #1

See

Tseq = Time taken by any process to complete sequentially on a single processor

Tseq = N * P

where N is no .of instruction and P is clock period

and here as nothing specifically given so time given in 20sec to execute on a single processor so i think that

Tseq= 20sec

Speed Up - Means how much new system is fast as compared to old system

So

Speed Up = time taken by single processor / time taken to run on a N proceesor system

Here n =4 as new system is 4 processor system

So speed Up= 20/5 = 4

Thank You '

Please if you find anything wrong first comment it and let me know before rating it otheriwse  it you like please rate it .